A railway truck of mass 20 tonnes, moving at 6.5 km/h is brought to rest by a buffer stop. The buffer exerts a force of 22.5 kN initially and this force increases uniformly by 60 kN for each 1 m compression of the buffer. Neglecting any loss of energy at impact, find the maximum compression of the buffer and the time required for the truck to be brought to rest.

1
Expert's answer
2021-03-24T19:39:52-0400
"K=W\\0.5mv^2=Fx+0.5kx^2\\0.5(20000)\frac{6.5^2}{3.6^2}=22500x+0.5(60000)x^2\\x=0.733\ m"